Sam, Choi, Mich, Pia, and Sheen are on a 6’ x 4’ rectangular balsa. They are having trouble in keeping the balsa balanced as they are unaware of their center of mass. Find their center of mass, given the following: Sheen is at 1 foot west of the balsa’s geometric center. She weighs 75 kg. 

1
Expert's answer
2021-10-28T13:49:37-0400

In general


"x=\frac{m_1x_1+m_2x_2+...+m_nx_n}{m_1+m_2+...+m_n}" and "y=\frac{m_1y_1+m_2y_2+...+m_ny_n}{m_1+m_2+...+m_n}" .


Data is specified for Sheen, but not for his friends.

Therefore, more data are needed to determine the center of mass.